if the temp. at noon is 78 degrees and at 5:00 it is 58 degrees how much did the temp. change each hour
Answer by checkley71(8403)

(78-58)/5=20/5=4 DROPPED AN AVERAGE OF 4 DEGREES EACH HOUR.